OpenStudy (anonymous):

Will give a medal for the correct answer... if sin theta = 12/13 and cos theta < 0, what is tan theta.

OpenStudy (anonymous):

A.13 B.13/12 C. -12/5 D. 12/5 E. - 5/12

OpenStudy (anonymous):

yttrium PLEASE HELP

OpenStudy (yttrium):

sin = opp/hyp cos = adj/hyp <0 Therefore, we have (-x,y) By using pythagorean theorem, we will get the third side which is the adjacent side of theta. So, cos(theta) = -5/13 Thereofore, by applying tan = sin/cosx. What will you get?

OpenStudy (yttrium):

@turtle123

OpenStudy (anonymous):

A.13 B.13/12 C. -12/5 D. 12/5 E. - 5/12

OpenStudy (anonymous):

I will give you a medal if you give me THE correct ANSWER...............

OpenStudy (debbieg):

@Turtle123 it is a violation of the site rules to ask for answers. OpenStudy values the Learning process - not the ‘Give you an answer’ process. Don’t post only answers - guide the asker to a solution. http://openstudy.com/code-of-conduct @Yttrium gave you everything you need to find the answer. Just read what he wrote, and ask questions if you don't understand something. This is a straight-forward problem and if you actually learn HOW to do it, you will be able to do a similar problem on your own next time, in less time than it takes to have someone help you with it here.

OpenStudy (debbieg):

And by the way, most of the people here who can help you LEARN care far more about THAT, than about getting a medal.